# Weyl Law and convergence in the classical limit for min-max nonlocal minimal surfaces

### v3

@@ Abstract
We study nonlocal minimal surfaces as a new approximation theory for the area functional, and more specifically in the context of Yau’s conjecture on the existence of minimal surfaces in closed three-dimensional manifolds.
In recent work, the author and collaborators constructed infinitely many nonlocal $s$-minimal hypersurfaces (via min-max methods) on any closed $n$-dimensional Riemannian manifold $M$, establishing a full analogue of Yau’s conjecture for $s\in(0,1)$.
The present article first proves a Weyl Law for the fractional perimeters of these hypersurfaces. The rest—and main part—of the article is devoted to obtaining effective estimates (as $s\to 1$) for min-max $s$-minimal surfaces in closed three-dimensional manifolds, including classical perimeter and mean curvature estimates, and eventually establishing their convergence to smooth, classical minimal surfaces. We recover in particular recent results on generic density and equidistribution of minimal surfaces, which are a strong form of Yau’s conjecture in this setting.
@@ 1 Introduction and main results
This gives a nonlocal analog of the Weyl-type Law for classical minimal hypersurfaces, which was conjectured by Gromov [35] in general codimension and proved by Liokumovich–Marques–Neves [40] in the case of minimal hypersurfaces obtained via the Almgren–Pitts min-max procedure. Soon after that, Gaspar–Guaraco [31] showed that the Weyl Law holds for limits (as $\varepsilon\to 0$) of min-max Allen–Cahn solutions as well, with a similar proof to the one in [40]. Interestingly, their Weyl Law can hold only for the limits obtained as $\varepsilon\to 0$; on the other hand, Theorem 1.1 holds already for any fixed $s\in(0,1)$. We emphasise here that $s$-minimal hypersurfaces are geometric objects on their own, with a canonical definition on Riemannian manifolds (see Section 3.1.1 and [17]), and moreover the parameter $s$ is “dimensionless”; in the Allen–Cahn case, $\varepsilon$ has “length units”, thus being an Allen–Cahn solution for a certain $\varepsilon$ is not a scaling invariant notion anymore, and solutions with larger energy (i.e. more transitions) can only exist when making $\varepsilon$ smaller.
@@
$\lim_{i\to\infty}(1-s_{i})\textnormal{Per}_{s_{i}}(E_{s_{i}})=\gamma_{n}\sum_{i=k}^{m}n_{k}\textnormal{Per}(\Sigma_{k})\,.$
@@
2 Overview of the proofs and further directions
@@
The proof of the compactness of $s$-minimal surfaces with bounded index occupies the rest (and main bulk) of our work. Despite the “clean” nature of most of the proofs and results, this requires a lengthier exposition. In fact, most of the results are completely new: As an example, our very first step is to obtain the (very challenging, yet deceptively simple-looking) formula for the second variation of the fractional perimeter of a domain on a manifold, which will require considering a smooth extension to the entire manifold of the normal vector of the domain, and will then take the form of a singular integral which will be shown to converge due to second order cancellations.
One of the new points we consider most interesting in our strategy is the eventual proof of global estimates for $s$-minimal surfaces with bounded index, including the uniform, quantitative $L^{1}$ decay of their classical mean curvatures as $s\to 1$ which holds even around points of singular convergence, and which will allows us to obtain the convergence in the varifold sense to a limit. Once the limit varifold is obtained, we will show its regularity outside of finitely many points; then, a careful argument using this partial regularity will allow us to show the convergence of the second (or any) variations of the approximating $s$-minimal surfaces to those of the limit, as well as the convergence of the fractional perimeters, and only then we will remove the singularities thanks to the stability properties inherited by the limit. We emphasise here that, to our knowledge, this is the first instance of a compactness theorem for $s$-minimal surfaces as $s\to 1$ without the presence of uniform curvature estimates; in particular, the convergence of critical points under exclusively an area bound assumption, as in [1] or [37], is an open and very interesting problem.
@@ 3.1.2 Fundamental properties
see for instance [10, Proposition 4.2]. The bound (23) then follows from applying this result to $v=u\circ\varphi$ and using Lemma 3.10 to compare the corresponding kernels. ∎
@@ 3.1.3 Monotonicity formula
Theorem 3.15 ([17, Theorem 2.25 and Proposition 3.2]).
@@
Theorem 3.16 (Monotonicity formula [17, Theorem 3.4]).
@@
We will now continue by proceeding as in [14, Proposition 7.1]. Let $\xi$ be a standard cutoff on $\widetilde{\mathcal{B}}_{4}^{+}\subset\mathbb{R}^{n+1}$, verifying that $\chi_{\widetilde{\mathcal{B}}_{2}^{+}\cap\{z=0\}}\leq\xi\leq\chi_{\widetilde{\mathcal{B}}_{4}^{+}}$. Define then $\widetilde{\eta}_{r}=\xi(\cdot/r)\circ\widetilde{\varphi}^{-1}$, extended by zero outside of $\widetilde{\varphi}(\widetilde{\mathcal{B}}_{4r}^{+})\subset\widetilde{M}$. Moreover, set $\eta_{r}(p)=\widetilde{\eta}_{r}(p,0):M\to\mathbb{R}$. Then, $\eta_{r}(p)$ acts as a cutoff function on $\varphi(\mathcal{B}_{4r})\subset M$ which verifies that $\chi_{\varphi(\mathcal{B}_{2r})}\leq\eta_{r}\leq\chi_{\varphi(\mathcal{B}_{4r})}$. We can then compute
@@
We now need essentially the converse of Step 1 to bound the RHS. This is [17, Lemma 3.15], or more precisely the chain of inequalities at the end of its proof, and it allows us to conclude that
@@ 3.2 Min-max procedure – The nonlocal volume spectrum
Critical points of (13) can be obtained using an equivariant min-max procedure, exploiting the $\mathbb{Z}_{2}$-symmetry of the functional $\mathcal{E}_{\varepsilon,s}$. In [18], a “nonlocal spectrum” is defined employing as a min-max family of sets with $\mathfrak{p}$ parameters the one consisting of odd images of the sphere $\mathbb{S}^{\mathfrak{p}}$ into $H^{s/2}(M)$. This family has a very simple definition and served for the purposes of that article, but others could have been employed in an essentially exchangeable way, including families which come from a topological index; see [18, Remark 3.1] where this is briefly discussed. To obtain a Weyl Law, one wants to restrict instead to families which arise from some topological index.
@@
We explain the slight modifications needed with respect to [18, Section 3.1] in Appendix A. ∎
@@ 4.1.2 Euclidean domain case
We will need the following simple result, corresponding to [40, Lemma 3.5]:
@@ 5.1 First and second variation formulas
We obtain for the first time a formula for the second variation of the fractional perimeter on a Riemannian manifold. The Euclidean formula, which was obtained in [25, Theorem 6.1], contains a difference of normal vectors based at different points, which does not have a Riemannian analogue. Moreover, a nonlocal correspondent of the Ricci curvature should implicitly appear. For these reasons, both the formula and its proof need fundamental modifications.
@@
Finally, arguing as in the proof of Proposition 3.11, which corresponds to [17, Proposition 3.9] (recall that $K^{\prime}_{\varepsilon}(p,q)=\frac{d}{dt}\big{|}_{t=0}K_{\varepsilon}(\psi_{\nu_{\partial E}}^{t}(p),\psi_{\nu_{\partial E}}^{t}(q))$), we see that we can bound
@@ 5.2 A first classical perimeter estimate for $s$-minimal surfaces with bounded index
The former is the only case we will need in this article. If, on the other hand, we did not assume (88) and wished to obtain (87) instead, we would need to continue the argument as in the proof of [18, Proposition 3.14]. First, we would use the interpolation result in (24) and the inequality $ab\leq\frac{a^{2}}{\delta}+\delta b^{2}$, valid for any $\delta>0$, to get
@@
for all $x\in\mathcal{B}_{3/4}$ and $r\leq 1/8$. Using then a covering lemma from [50], see [18, Lemma 3.18] for the exact statement, we would then conclude that $\text{Per}(E,\varphi(\mathcal{B}_{1/2}))\leq\frac{C}{1-s}$ as desired.
@@
The proof of Theorem 5.4 in the finite index case will need the following covering-type lemma, which was devised in [18] to exhibit estimates in the finite Morse index case by iteratively reducing to the almost-stable case. It was inspired by the proof of [26, Proposition 2.6]. Essentially, it says that if every time we take $(m+1)$ cubes which are sufficiently far from each other we are able to prove an estimate of power-type on at least one of them, then the estimate actually holds everywhere. The proof consists in subdividing a cube at scale 1 into small subcubes, adding the estimate on the small subcubes (except for a “bad” set, which will consist of at most $m$ subcubes and their close neighbours), and then iterating the procedure dyadically by subdividing the “bad” subcubes once again into smaller subcubes and starting the argument again.
Lemma 5.11 ([18, Lemma 3.19]).
@@ 5.3.2 Uniform $C^{2,\alpha}$ estimates for almost-stable graphs in dimension $n=3$
It is well known that the classical mean curvature can be written, for a graph in coordinates, as an operator depending on the graph and the coefficients of the metric. Moreover, the difference of the parametrising functions of two parallel graphs satisfies an elliptic equation. The precise statement is that, for $x^{\prime}\in B^{\prime}_{3/4}$ and $v:=g_{i+1}-g_{i}$, we have that (see for example [22, page 237])
@@
On the other hand, by the Harnack inequality for uniformly elliptic operators, see [34, Theorems 8.17 and 8.18], we have that
@@
We now argue as in [20, Proposition 3.5]. Still in dimension $n=3$ (and thus $n^{\prime}=2$), assume that $(g_{i+1}-g_{i})(x_{\circ}^{\prime})=\delta\sqrt{\sigma}$ for some $x^{\prime}_{\circ}\in B^{\prime}_{1/2}$ and let us prove a lower bound for $\delta$.
@@
On the other hand, the $\varphi^{-1}(A_{j})$ are sets with uniformly $C^{2}$ (graphical) boundaries and which are transversal to $\partial\Omega_{0}$ in a quantified way thanks to their graphical structure. We can then apply [16, Lemma 11] separately to each of the $\varphi^{-1}(A_{j})$, obtaining that
@@ 5.3.3 Uniform $C^{2,\alpha}$ regularity for almost-stable surfaces in dimension $n=3$
Step 1. Let us show first that necessarily $s_{k}\to 1$. Indeed, suppose (up to extracting a subsequence) that $s_{k}\to s\in[s_{*},1)$. Then, thanks to (uniform, since $s_{k}$ is bounded away from 1) layer separation estimates —here we can even use some rough layer separation estimate as in Lemma 5.14— and the standard $C^{3}$ regularity of fractional minimal surfaces (e.g. arguing as in the last part of the proof of Corollary 5.25), by Ascoli–Arzelà we would find that the $\partial\widetilde{F}_{k}$ converge locally in $C^{2}$ (as submanifolds of $\mathbb{R}^{3}$) to an embedded hypersurface $\partial\widetilde{F}\subset\mathbb{R}^{3}$ of class $C^{2}$. Moreover, by Remark 5.9 the $\partial E_{k}$ are $\Lambda/R_{k}^{n+s_{k}-2}$-almost stable in $(M_{k},R_{k}^{2}g_{k})$, which since $R_{k}\to\infty$ gives immediately that $\partial\widetilde{F}$ must be a stable $s$-minimal surface. See e.g. Step 2 of the proof of [13, Proposition 6.1] for a similar argument on how stability passes to the limit in the present case. Hence by the Bernstein–type result for $C^{2}$ stable $s$-minimal surfaces in [13, Corollary 2.12], $\widetilde{F}$ must be a half-space. On the other hand, passing to the limit the equation $|{\rm I\negthinspace I}_{\partial\widetilde{F}_{k}}(0)|=1$ we obtain, thanks to the $C^{2}$ convergence, that $|{\rm I\negthinspace I}_{\partial\widetilde{F}}(0)|=1$, which is a contradiction.
@@
The proof is as the one for (76), defining the single variable functions $f_{k}(t)=(1-s_{k})\textnormal{Per}_{s_{k}}(\psi_{X_{k}}^{t}(E_{k});\widetilde{\varphi}_{k}(\mathcal{B}_{R}))$ and $f(t)=\gamma_{n}\|(\psi_{X}^{t})_{\#}V\|_{\mathcal{B}_{R}}$, showing that $f_{k}(t)\to f(t)$ as $k\to\infty$ thanks to Lemma 5.26 as in (137), and applying Lemma 3.12 to see that the functions $\frac{d^{\ell}}{dt^{\ell}}f_{k}(t)$ are locally uniformly bounded. See [18, Lemma 4.4] for full details. ∎

### v4

@@ Abstract
We study nonlocal minimal surfaces as a new approximation theory for the area functional, and more specifically in the context of Yau’s conjecture on the existence of minimal surfaces in closed three-dimensional manifolds. This programme offers an alternative to the Almgren–Pitts and Allen–Cahn approaches, with advantageous features both from the existence and regularity viewpoints.
We build on recent work in which the author and collaborators constructed infinitely many nonlocal $s$-minimal hypersurfaces (via min-max methods) on any closed $n$-dimensional Riemannian manifold, establishing a full analogue of Yau’s conjecture for $s\in(0,1)$.
The present article first proves a Weyl-type Law for the fractional perimeters of these hypersurfaces. The rest—and main part—of the article is devoted to obtaining uniform estimates (in the classical limit $s\to 1$) for min-max $s$-minimal surfaces in closed three-manifolds, eventually establishing their convergence to smooth, classical minimal surfaces. We recover in particular recent results on existence, generic density and equidistribution of minimal surfaces, which are a strong form of Yau’s conjecture in this setting.
@@ 1 Introduction and main results
This gives a nonlocal analogue of the Weyl-type Law for classical minimal hypersurfaces, which was conjectured by Gromov [37] in general codimension and proved by Liokumovich–Marques–Neves [42] in the case of minimal hypersurfaces obtained via the Almgren–Pitts min-max procedure. Soon after that, Gaspar–Guaraco [33] showed that the Weyl Law holds for limits (as $\varepsilon\to 0$) of min-max Allen–Cahn solutions as well, with a similar proof to the one in [42]. Interestingly, their Weyl Law can hold only for the limits obtained as $\varepsilon\to 0$; on the other hand, Theorem 1.1 holds already for any fixed $s\in(0,1)$. We emphasise here that $s$-minimal hypersurfaces are geometric objects on their own, with a canonical definition on Riemannian manifolds (see Section 3.1.1 and [18]), and moreover the parameter $s$ is “dimensionless”; in the Allen–Cahn case, $\varepsilon$ has “length units”, thus being an Allen–Cahn solution for a certain $\varepsilon$ is not a scaling invariant notion anymore, and solutions with larger energy (i.e. more transitions) can only exist when making $\varepsilon$ smaller.
@@
$\lim_{i\to\infty}(1-s_{i})\textnormal{Per}_{s_{i}}(E_{s_{i}})=\gamma_{n}\sum_{k=1}^{m}n_{k}\textnormal{Per}(\Sigma_{k})\,.$
@@
2 Overview of the article and further directions
@@
The proof of the compactness of $s$-minimal surfaces with bounded index occupies the rest (and main bulk) of our work. Most of the results are of a remarkably “clean” nature, providing—together with [19]—a self-contained and robust (as $s\to 1$) regularity theory for min-max $s$-minimal surfaces. Combined with the elementary existence theory in [19], this showcases the nonlocal approximation as an attractive alternative to the Almgren–Pitts and Allen–Cahn approaches. For comparison, the Allen–Cahn approach offers a simple existence theory too, but at the expense of a very hard regularity theory:
  • In general dimension, the only available results are via the deep and powerful regularity theory in [57] for stable integral varifolds. This method presents moreover important difficulties, such as a potential loss of geometric information in the prescribed mean curvature [8] or manifold-with-boundary [45] cases.
  • In dimension $n=3$, a stronger, alternative regularity theory is present, developed in the (similarly deep) breakthrough articles [55, 56, 23]. An extension to higher dimensions depends on the longstanding open problem of classification of stable solutions (with bounded energy density) to Allen–Cahn, which so far has only been achieved for $n=4$ in the very recent (independent) article [31].
Now, despite the “clean” nature of the nonlocal approach, the regularity and compactness part of the article will require a lengthier exposition. The main reason is that, in fact, we need to build most of the theory from scratch: As an example, our very first step is to obtain the (challenging, yet deceptively simple-looking) formula for the second variation of the fractional perimeter of a domain on a manifold, which will require considering a smooth extension to the entire manifold of the normal vector of the domain, and will then take the form of a singular integral which will be shown to converge due to second order cancellations.
A point we consider most interesting in our strategy is the eventual proof of global estimates for $s$-minimal surfaces with bounded index, including the uniform, quantitative $L^{1}$ decay of their classical mean curvatures as $s\to 1$ which holds even around points of singular convergence, and which will allows us to obtain the convergence in the varifold sense to a limit. Once the limit varifold is obtained, we will show its regularity outside of finitely many points; then, a careful argument using this partial regularity will allow us to show the convergence of the second (or any) variations of the approximating $s$-minimal surfaces to those of the limit, as well as the convergence of the fractional perimeters, and only then we will remove the singularities thanks to the stability properties inherited by the limit. We emphasise here that, to our knowledge, this is the first instance of a compactness theorem for $s$-minimal surfaces as $s\to 1$ without the presence of uniform curvature estimates; in particular, the convergence of critical points under exclusively an area bound assumption, as in [1] or [39], is an open and very interesting problem.
@@ 3.1.2 Fundamental properties
see for instance [11, Proposition 4.2]. The bound (23) then follows from applying this result to $v=u\circ\varphi$ and using Lemma 3.10 to compare the corresponding kernels. ∎
@@ 3.1.3 Monotonicity formula
Theorem 3.15 ([18, Theorem 2.25 and Proposition 3.2]).
@@
Theorem 3.16 (Monotonicity formula [18, Theorem 3.4]).
@@
We will now continue by proceeding as in [15, Proposition 7.1]. Let $\xi$ be a standard cutoff on $\widetilde{\mathcal{B}}_{4}^{+}\subset\mathbb{R}^{n+1}$, verifying that $\chi_{\widetilde{\mathcal{B}}_{2}^{+}\cap\{z=0\}}\leq\xi\leq\chi_{\widetilde{\mathcal{B}}_{4}^{+}}$. Define then $\widetilde{\eta}_{r}=\xi(\cdot/r)\circ\widetilde{\varphi}^{-1}$, extended by zero outside of $\widetilde{\varphi}(\widetilde{\mathcal{B}}_{4r}^{+})\subset\widetilde{M}$. Moreover, set $\eta_{r}(p)=\widetilde{\eta}_{r}(p,0):M\to\mathbb{R}$. Then, $\eta_{r}(p)$ acts as a cutoff function on $\varphi(\mathcal{B}_{4r})\subset M$ which verifies that $\chi_{\varphi(\mathcal{B}_{2r})}\leq\eta_{r}\leq\chi_{\varphi(\mathcal{B}_{4r})}$. We can then compute
@@
We now need essentially the converse of Step 1 to bound the RHS. This is [18, Lemma 3.15], or more precisely the chain of inequalities at the end of its proof, and it allows us to conclude that
@@ 3.2 Min-max procedure – The nonlocal volume spectrum
Critical points of (13) can be obtained using an equivariant min-max procedure, exploiting the $\mathbb{Z}_{2}$-symmetry of the functional $\mathcal{E}_{\varepsilon,s}$. In [19], a “nonlocal spectrum” is defined employing as a min-max family of sets with $\mathfrak{p}$ parameters the one consisting of odd images of the sphere $\mathbb{S}^{\mathfrak{p}}$ into $H^{s/2}(M)$. This family has a very simple definition and served for the purposes of that article, but others could have been employed in an essentially exchangeable way, including families which come from a topological index; see [19, Remark 3.1] where this is briefly discussed. To obtain a Weyl Law, one wants to restrict instead to families which arise from some topological index.
@@
We explain the slight modifications needed with respect to [19, Section 3.1] in Appendix A. ∎
@@ 4.1.2 Euclidean domain case
We will need the following simple result, corresponding to [42, Lemma 3.5]:
@@ 5.1 First and second variation formulas
We obtain for the first time a formula for the second variation of the fractional perimeter on a Riemannian manifold. The Euclidean formula, which was obtained in [26, Theorem 6.1], contains a difference of normal vectors based at different points, which does not have a Riemannian analogue. Moreover, a nonlocal correspondent of the Ricci curvature should implicitly appear. For these reasons, both the formula and its proof need fundamental modifications.
@@
Finally, arguing as in the proof of Proposition 3.11, which corresponds to [18, Proposition 3.9] (recall that $K^{\prime}_{\varepsilon}(p,q)=\frac{d}{dt}\big|_{t=0}K_{\varepsilon}(\psi_{\nu_{\partial E}}^{t}(p),\psi_{\nu_{\partial E}}^{t}(q))$), we see that we can bound
@@ 5.2 A first classical perimeter estimate for $s$-minimal surfaces with bounded index
The former is the only case we will need in this article. If, on the other hand, we did not assume (88) and wished to obtain (87) instead, we would need to continue the argument as in the proof of [19, Proposition 3.14]. First, we would use the interpolation result in (24) and the inequality $ab\leq\frac{a^{2}}{\delta}+\delta b^{2}$, valid for any $\delta>0$, to get
@@
for all $x\in\mathcal{B}_{3/4}$ and $r\leq 1/8$. Using then a covering lemma from [53], see [19, Lemma 3.18] for the exact statement, we would then conclude that $\text{Per}(E,\varphi(\mathcal{B}_{1/2}))\leq\frac{C}{1-s}$ as desired.
@@
The proof of Theorem 5.4 in the finite index case will need the following covering-type lemma, which was devised in [19] to exhibit estimates in the finite Morse index case by iteratively reducing to the almost-stable case. It was inspired by the proof of [27, Proposition 2.6]. Essentially, it says that if every time we take $(m+1)$ cubes which are sufficiently far from each other we are able to prove an estimate of power-type on at least one of them, then the estimate actually holds everywhere. The proof consists in subdividing a cube at scale 1 into small subcubes, adding the estimate on the small subcubes (except for a “bad” set, which will consist of at most $m$ subcubes and their close neighbours), and then iterating the procedure dyadically by subdividing the “bad” subcubes once again into smaller subcubes and starting the argument again.
Lemma 5.11 ([19, Lemma 3.19]).
@@ 5.3.1 $L^{-2}$ separation estimate for almost-stable graphs
${\mbox{satisfying}}\quad g_{1}<g_{2}<\cdots<g_{N}\qquad\mbox{and}\qquad\|\nabla g_{i}\|_{L^{\infty}(\mathcal{B}_{1}^{\prime})}+\|D^{2}g_{i}\|_{L^{\infty}(\mathcal{B}_{1}^{\prime})}<\delta\,.$
@@ 5.3.2 Uniform $C^{2,\alpha}$ estimates for almost-stable graphs in dimension $n=3$
It is well known that the classical mean curvature can be written, for a graph in coordinates, as an operator depending on the graph and the coefficients of the metric. Moreover, the difference of the parametrising functions of two parallel graphs satisfies an elliptic equation. The precise statement is that, for $x^{\prime}\in B^{\prime}_{3/4}$ and $v:=g_{i+1}-g_{i}$, we have that (see for example [23, page 237])
@@
On the other hand, by the Harnack inequality for uniformly elliptic operators, see [36, Theorems 8.17 and 8.18], we have that
@@
We now argue as in [21, Proposition 3.5]. Still in dimension $n=3$ (and thus $n^{\prime}=2$), assume that $(g_{i+1}-g_{i})(x_{\circ}^{\prime})=\delta\sqrt{\sigma}$ for some $x^{\prime}_{\circ}\in B^{\prime}_{1/2}$ and let us prove a lower bound for $\delta$.
@@
On the other hand, the $\varphi^{-1}(A_{j})$ are sets with uniformly $C^{2}$ (graphical) boundaries and which are transversal to $\partial\Omega_{0}$ in a quantified way thanks to their graphical structure. We can then apply [17, Lemma 11] separately to each of the $\varphi^{-1}(A_{j})$, obtaining that
@@ 5.3.3 Uniform $C^{2,\alpha}$ regularity for almost-stable surfaces in dimension $n=3$
Step 1. Let us show first that necessarily $s_{k}\to 1$. Indeed, suppose (up to extracting a subsequence) that $s_{k}\to s\in[s_{*},1)$. Then, thanks to (uniform, since $s_{k}$ is bounded away from 1) layer separation estimates —here we can even use some rough layer separation estimate as in Lemma 5.14— and the standard $C^{3}$ regularity of fractional minimal surfaces (e.g. arguing as in the last part of the proof of Corollary 5.25), by Ascoli–Arzelà we would find that the $\partial\widetilde{F}_{k}$ converge locally in $C^{2}$ (as submanifolds of $\mathbb{R}^{3}$) to an embedded hypersurface $\partial\widetilde{F}\subset\mathbb{R}^{3}$ of class $C^{2}$. Moreover, by Remark 5.9 the $\partial E_{k}$ are $\Lambda/R_{k}^{n+s_{k}-2}$-almost stable in $(M_{k},R_{k}^{2}g_{k})$, which since $R_{k}\to\infty$ gives immediately that $\partial\widetilde{F}$ must be a stable $s$-minimal surface. See e.g. Step 2 of the proof of [14, Proposition 6.1] for a similar argument on how stability passes to the limit in the present case. Hence by the Bernstein–type result for $C^{2}$ stable $s$-minimal surfaces in [14, Corollary 2.12], $\widetilde{F}$ must be a half-space. On the other hand, passing to the limit the equation $|{\rm I\negthinspace I}_{\partial\widetilde{F}_{k}}(0)|=1$ we obtain, thanks to the $C^{2}$ convergence, that $|{\rm I\negthinspace I}_{\partial\widetilde{F}}(0)|=1$, which is a contradiction.
@@
The proof is as the one for (76), defining the single variable functions $f_{k}(t)=(1-s_{k})\textnormal{Per}_{s_{k}}(\psi_{X_{k}}^{t}(E_{k});\widetilde{\varphi}_{k}(\mathcal{B}_{R}))$ and $f(t)=\gamma_{n}\|(\psi_{X}^{t})_{\#}V\|_{\mathcal{B}_{R}}$, showing that $f_{k}(t)\to f(t)$ as $k\to\infty$ thanks to Lemma 5.26 as in (137), and applying Lemma 3.12 to see that the functions $\frac{d^{\ell}}{dt^{\ell}}f_{k}(t)$ are locally uniformly bounded. See [19, Lemma 4.4] for full details. ∎